help! will rate!
A 0.25 M NaA solution has a pH of 9.25 at 25°C. Calculate the pK value of HA.
Answer #1

ANSWEX! bien that, [Nan] = [] = 0.25M pH=9.95 Hode, we are given the conjugate base of weak acid Hm. A + H₂O ---) OH + HA 0 0= 1784105m So at ewlibrium, [ok-3= Y = 1.78X10-50% At equilibrium, .:. [^] = 0.25-1.78X16-57 = 0,25 [HA] = X =17881-5 go, bag